# Near-quantum limited axion dark matter search with the ORGAN experiment around 26 $\mu$eV

### v2

@@ Abstract
We have reported the results from the first near-quantum limited phase of The ORGAN Experiment, operated at millikelvin temperatures and with a commercially available JPA. It is the most sensitive probe on the axion-photon coupling strength in the mass range between $25.45-26.27\,\mu\text{eV}$, excluding $|g_{a\gamma\gamma}|\gtrsim 2.8\times 10^{-13}$ at a 95% confidence level.
@@ Acknowledgements.
Supplementary Material for: Near-quantum limited axion dark matter search with the ORGAN experiment around 26 $\mu$eV
Aaron P. Quiskamp,1,* Graeme R. Flower,1 Steven Samuels,1 Ben T. McAllister,1,2$\dagger$ Paul Altin,3 Eugene N. Ivanov,1 Maxim Goryachev,1 Michael E. Tobar1$\ddagger$
1Quantum Technologies and Dark Matter Laboratory,
Department of Physics, University of Western Australia,
35 Stirling Highway, Crawley WA 6009, Australia.
2ARC Centre of Excellence for Dark Matter Particle Physics,
Swinburne University of Technology, John St, Hawthorn VIC 3122, Australia
3ARC Centre of Excellence For Engineered Quantum Systems,
The Australian National University, Canberra ACT 2600 Australia
I JPA characterisation
@@
Supplementing the main text, some further details as to the characterisation and use of the JPA are provided. As discussed in the main text, the first stage amplifier used in the experimental run was a commercially available JPA from Raytheon [37]. Due to the non-standard nature of the input coupling (a broadband impedance transformer to the junction resonator), the JPA is operated in a different way to typical capacitively coupled JPAs. This has the significant benefit of producing broadband gain beyond the gain bandwidth product - but does complicate tune up, as optimal gain and noise performance is often achieved with pump tones significantly detuned from resonance.
@@
The junction is in a SQUID configuration, with flux through the loop driven by the pump port, whilst DC current applied to the flux port tunes the JPA resonance. Due to the resonance being highly over-coupled to a 50 $\Omega$ load, it is characterized via the normalised reflected phase as shown in Fig. 4.
As the device is cooled from above the transition temperature of aluminium (approximately 1.2K) to below, it traps a random quantity of flux in the SQUID loop resulting in small changes to the tuning parameters of the device compared to previous cools, requiring re-characterisation each time the device makes that transition. Unfortunately, due to magnet quenches that occurred during the run resulting in a rapid warm up of the dilution refrigerator, the JPA required re-characterisation. A list of JPA pump and bias parameters used over particular frequencies is presented in table 2. Additionally, gains associated with these pump and bias parameters are noted in Fig. 4.
II Data Analysis
As discussed in the main text, we follow the HAYSTAC analysis procedure [41]. The analysis window is restricted to an IF bandwidth of $\sim 3\,\text{MHz}$, and the cropped spectra are subsequently normalized using a Savitsky-Golay (SG) filter with a 101-point window and a polynomial degree of 4. The SG filter removes the slowly varying baseline from the raw spectra whilst preserving narrow spectral features on the scale of the axion width $\Delta\nu_{a}\sim 6.25\,\text{kHz}$. Overlapping power spectra are then vertically combined using a maximum likelihood weighted sum of contributing spectra to maximize the SNR. The weights for this sum are determined according to the expected axion power in the $j$th IF bin of the $i$th spectrum and the bin-dependent $T^{\text{S}}_{ij}$ value.
The “grand power spectrum” is constructed by optimally filtering sets of 9 consecutive bins with the expected Maxwellian lineshape in the boosted lab frame, retaining $\sim 90\%$ of the axion signal [41]. The standard deviation for the $k$th grand spectrum bin is then the quadrature weighted sum of contributing combined spectra standard deviations, denoted as $\sigma^{g}_{k}$. The grand power spectra are normalized to $\sigma^{g}_{k}$ and shown histogrammed in panel (a) of Fig. 5. The width of this distribution is reduced to $\xi^{g}=0.90$ due to the negative correlations induced by the SG filter among neighbouring bins. The SG filter is also known to attenuate potential axion signals, and is a function of the choice of filter parameters. This effect is simulated 1,000,000 times by injecting synthetic axions into the ORGAN-Q analysis pipeline and comparing it with unfiltered data [41, 23]. The result is shown in panel (b) of Fig. 5, where the ratio of the two means $\eta_{\text{SG}}=6.94/9.61=0.72$ gives the SG-filter-induced attenuation. This factor is reduced by $\xi^{g}$ to give $\eta_{\text{SNR}}=\eta_{\text{SG}}/\xi^{g}=0.80$, which represents the total reduction in SNR of each grand bin to a given axion signal [41]. To validate the simulation, we also compute $\eta_{\text{SG}}$ analytically based on the recent work in [50], and find $\eta_{\text{SG}}=0.73$.
We chose to forgo rescans due to multiple magnet quenches during the data-taking run, which significantly extended the experimental time. Bin-by-bin exclusion limits on axion-photon coupling are placed based on the maximum normalized power excess of $6.68\sigma$, which corresponds to an SNR target of $8.33\sigma$ at 95% confidence level [41]. Exclusion limits are then calculated based on that and presented in the main text.

### v3

@@ III Conclusion
We have reported the results from the first near-quantum-limited phase of the ORGAN experiment, operated at millikelvin temperatures and with a commercially available JPA. It is the most sensitive probe on the axion-photon coupling strength in the mass range between $25.45-26.27\,\mu\text{eV}$, excluding $|g_{a\gamma\gamma}|\gtrsim 2.8\times 10^{-13}$ at a 95% confidence level.
@@ Acknowledgements.
Appendix A: JPA characterisation
@@
Further details as to the characterisation and use of the JPA are provided. As discussed in the above, the first stage amplifier used in the experimental run was a commercially available JPA from Raytheon [38]. Due to the nonstandard nature of the input coupling (a broadband impedance transformer to the junction resonator), the JPA is operated in a different way to typical capacitively coupled JPAs. This has the significant benefit of producing broadband gain beyond the gain bandwidth product – but does complicate optimisation, as best gain and noise performance is often achieved with pump tones significantly detuned from resonance.
@@
The junction is in a SQUID configuration, with flux through the loop driven by the pump port, while DC current applied to the flux port tunes the JPA resonance. Due to the resonance being highly overcoupled to a 50 $\Omega$ load, it is characterized using the normalised reflected phase as shown in Fig. 5.
As the device is cooled from above the transition temperature of aluminium (approximately 1.2 K) to below, it traps a random quantity of flux in the SQUID loop resulting in small changes to the tuning parameters of the device compared to previous cools, requiring recharacterisation each time the device makes that transition. Unfortunately, due to magnet quenches that occurred during the run resulting in a rapid warm up of the dilution refrigerator, the JPA required recharacterisation. A list of JPA pump and bias parameters used over particular frequencies is presented in Table 2. Additionally, gains associated with these pump and bias parameters are noted in Fig. 5.
Appendix B: Uncertainty Analysis
The systematic uncertainties relevant to the experimental run are summarised in Table I of the main text. The uncertainty in the frequency-dependent form factor $C$ is assumed to be within the range reported for similar experiments, typically $3\%$ to $5\%$ [43, 15], from which we adopt the upper limit of $5\%$. The uncertainty in the frequency-dependent total system noise temperature $T_{\text{S}}$ is taken to be the standard deviation of its measured values. This approach captures the variability in $T_{\text{S}}$ across the frequency range of interest, reflecting the fluctuations in the measured SNRI and $T\mathrm{{}^{Y}_{HEMT}}$. We find the uncertainty in $T_{\text{S}}$ to be $\sim 11.6\%$. The parameters $Q_{L}$, $\beta$, and $B_{0}$ were measured with high precision and contribute a negligible uncertainty, $\leq 1\%$
Considering only the dominant sources of uncertainty, the fractional uncertainty in the axion-photon coupling, $\delta g_{a\gamma\gamma}$ is given by
$\frac{\delta g_{a\gamma\gamma}}{g_{a\gamma\gamma\gamma}}\approx\sqrt{\left(\frac{1}{2}\frac{\delta T_{\mathrm{S}}}{T_{\mathrm{S}}}\right)^{2}+\left(\frac{1}{2}\frac{\delta C}{C}\right)^{2}}\approx 6.3\%$ (8)